Three times the larger numbers is equal to four times the smaller. The sum of the numbers is 21. Find the numbers.
Answer by uma(370)

Let the larger number = x and the smaller be y.
Then we have x + y = 21 -------------(1)
Also given 3 times the larger = 4 times the smaller
==> 3x = 4y
==> 3x - 4y = 0
4x + 4y = 84 [ 4 times (1)]
Adding the above 2 equations, 7x = 84
==> x = 84/7
==> x = 12
So 3x = 36
==> 4y = 36
==> y = 36/4
==> y = 9
Thus the bigger number is 12 and the smaller number is 9.
good luck!!!